• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name eschulte

    Log entry time 16:09:22 on June 16, 2007

    Entry number 205642

    keyword=Day Shift Log

    08:00 Begin shift finishing kinematics 3, He target, radiator in, 30uA.

    08:15 MCC called to set up the plan for the pass change to begin at 08:30.

    08:30 MCC called to let us know that they are getting ready for the
    energy change. Changing the target to empty target and removing the
    radiator. Requested a controlled access so that Alex can check the
    trigger scintillators.

    08:45 HRS-L Q3 started drifting a bit. This doesn't really effect
    anything at the moment, since we're under a pass change. ARM arrives to
    let Alex and Richard make their access. The hall will be swept and
    dropped to controlled access.

    10:45 Lost communication with target and rebooted IOC. Called J. Larose
    (target on-call) to ask a few questions following reboot.

    11:25 After talking to J. Larose we've figured out that the IOC rebooted
    fine, Loop 3 is idle, and that the read-out for the fan for loop 1 on the
    STRIP CHART (not the target loop panels) isn't reading out properly.
    Alex is still in the hall.

    12:05 Begin setup for next kinematics point: Kinematics 4
    HRS-L Angle: 44.11 deg, HRS-L Momentum: 2.7818 GeV/c
    HRS-R Angle: 44.11 deg, HRS-R Momentum: 2.7818 GeV/c

    The angles were set without incident:

    HRS-L: 44.5 deg, -36.5 Vernier
    HRS-R: 44.0 deg, -15.0 Vernier

    12:26 Set momentum to 2.7818 GeV/c for both spectrometers after Q2/3
    cycle on each. So far they are setting just fine.

    13:12 MCC called to ask if we still needed the Controlled Access. I told
    them no and they are bringing the hall back to Beam Permit.

    15:14 MCC called to let us know they are going to send some tune beam to
    the hall.

    15:30 Other than a tedious alarm from the HRS-L Q3 drifting low all is
    quiet.

    15:37 Notice that the HRS-L Q2 Helium level is at 98%. Reset itself.
    Trying cycling of Q3 (HRS-L) to see if that will stop the annoying alarm.

    15:46 Cycling Q3 seems to have settled the drift a bit, maybe. The
    Dipole unsettled itself during resetting of Q3 (by momentum). It's
    settling down.

    15:52 MCC called to let us know that they were going to send low current
    CW beam to the hall. We still have empty target, no radiator.

    16:00 iocha4 lost communication and needed a reboot. Can't seem to
    access hareboot2, the ups with this ioc connected. Passed this on to the
    next shift leader.